Identify the proportions that are true.
A) 1/3=4/12
B) 1/2=6/12
C) 5/7=25/49
D) 5/6=11/12
E) none of these

Answer by Alan3354(69443)

Some reason you can't do that?
---------------
If a/b = c/d then a*d = b*c